OpenStudy (anonymous):

!!!!!!!!!!!!!!!(Quick help please) Write equation in slope intercept form. Passing through (-5,9) and perpendicular to the line whose equation is y+1=5/6(x-7). (Since this is in point-slope form how do I make it into slope-intercept, do I just add the point (-5,9) to the y+1 (y+1-5) and (x-7-9)? )

OpenStudy (triciaal):

given y+1=5/6(x-7). slope = 5/6 = m slope of perpendicular = -1/m = -1/(5/6) = -6/5 passing through (-5, 9) y-9= -6/5(x - -5) = -6/5(x + 5) add 9 to each side to change to slope intercept form, y = m x + b y = -6/5x + (9 - 6) y = -6/5 x + 3

OpenStudy (triciaal):

any questions?